OpenStudy (anonymous):

i have question on my equation A= 7500(1.0000273972)

OpenStudy (anonymous):

so i already started A= 7500(1+0.01/365)^(365*10) A= 7500(1.0000273972)^(3650) A= 7500(1.0000273972)

OpenStudy (anonymous):

what's next ?? I'm completely stuck

OpenStudy (anonymous):

I have an example on my lesson and it's A=5000(1+0.0075/365)^365*10 A=5000(1.0000205479)^3650 A=5000(1.0778833) A=5389.42

OpenStudy (anonymous):

i understand this but in the example equation from step 2 to 3 and 4 what where the steps to get to that

OpenStudy (cwrw238):

you can only do the third line by using a scientific or graphical calculator enter 1.0000273972 then ^ then 3650 Press EXE and it will give you the result The last part is multiplying by 7500

OpenStudy (anonymous):

ok so i would need to first do 1.0000273972^3650 then multiply 7500

OpenStudy (anonymous):

hey when i do that on my lesson example i get 1.07788314249

OpenStudy (anonymous):

not 1.0778833

OpenStudy (anonymous):

i got 1.10516916104 @cwrw238

OpenStudy (cwrw238):

looks good

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

is there another way of doing this @cwrw238, my teacher gave me an F